OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (anonymous):

y-intercept: y = -2(0) - 3 y = -3 x-intercept: 0 = -2x - 3 -2x = 3 x = -(3/2)